# Adaptive Human Trajectory Prediction via Latent Corridors

### v1

@@ Abstract
Human trajectory prediction is typically posed as a zero-shot generalization problem: a predictor is learnt on a dataset of human motion in training scenes, and then deployed on unseen test scenes. While this paradigm has yielded tremendous progress, it fundamentally assumes that trends in human behavior within the deployment scene are constant over time. As such, current prediction models are unable to adapt to scene-specific transient human behaviors, such as crowds temporarily gathering to see buskers, pedestrians hurrying through the rain and avoiding puddles, or a protest breaking out. We formalize the problem of scene-specific adaptive trajectory prediction and propose a new adaptation approach inspired by prompt tuning called latent corridors. By augmenting the input of any pre-trained human trajectory predictor with learnable image prompts, the predictor can improve in the deployment scene by inferring trends from extremely small amounts of new data (e.g., 2 humans observed for 30 seconds). With less than $0.1\%$ additional model parameters, we see up to $23.9\%$ ADE improvement in MOTSynth simulated data and $16.4\%$ ADE in MOT and Wildtrack real pedestrian data. Qualitatively, we observe that latent corridors imbue predictors with an awareness of scene geometry and scene-specific human behaviors that non-adaptive predictors struggle to capture.
Figure 1: Adaptive trajectory prediction. (left) Given a history of human behavior (shown in black), the pre-trained predictor $\mathcal{P}$ is unable to understand scene-specific behavior trends, like people entering a subterranean subway entrance (bottom row). (right) When adapting, the number of people and amount of time determine the total number of trajectories observed, and we denote this time-dependent quantity human-seconds. Here, the three columns correspond to our method trained for a very small (left), medium (middle), and large amount of human-seconds. Our adaptive latent corridors approach enables $\mathcal{P}$ to quickly learn scene-specific trends, improving predictions with even small amounts of data, and closing the gap between the ground-truth (green) and predicted behavior (orange). For example, in the middle row, we see that $\mathcal{P}$ predicts the person will move towards the camera, but as our method sees more human-seconds of data, it adapts to the trend that people in this plaza scene tend to avoid the center of the plaza and instead move diagonally across it.
@@ 1 Introduction
Human motion prediction is a fundamental skill for intelligent systems to effectively navigate the world, assist end-users, and visually survey a scene. To date, learning-based human trajectory prediction has been extensively studied, making huge strides in predicting multimodal future behavior [27], modeling multi-agent social interactions [2, 17, 35], and accounting for scene context [49, 28].
@@
However, a fundamental assumption underlies the current trajectory prediction paradigm: predictors are assumed to be deployed within an unchanging context. Nevertheless, in the real world, the context—and thus patterns of human behavior—will inevitably change over time, as observed in [44]. For example, data from a security camera will capture a new subway entrance being built, and people descending and exiting in new patterns (bottom row, Fig. 1). Humans rushing to work during the day will largely ignore each other while nighttime party-goers will walk in cliques. Even if the trajectory predictor observes data captured from exactly the same location, the human behavior patterns can change in minutes or hours. When faced with such changes, the performance of existing trajectory predictors immediately degrades.
@@
Here, we study how to efficiently adapt pre-trained trajectory predictors to observed human behavior within a deployment scene (right, Fig. 1). We focus on adaptation that can be done in a data-efficient way and can preserve any useful structure within the base predictor in the new context (e.g., even if a new subway entrance lets people descend into the ground, other people may still move towards the existing above-ground coffee shops). Our technical approach takes inspiration from the recently popularized paradigm of prompt-tuning in large language models [23], but instantiates this idea within the adaptive trajectory prediction problem. Specifically, we augment the input to a human trajectory predictor with a set of latent image prompts, one per each instance of a scene we wish to adapt to. We leave all or most of the predictor frozen, and tune each input prompt with the same trajectory prediction loss that is used to train the base predictor. We find that even extremely small amounts of new human trajectories (e.g., $<1$ minute of human seconds, corresponding to 2 humans observed over 30 seconds) can leak sufficient information about how humans interact with the scene (e.g., new subway entrance) or with each other (e.g., nighttime social cliques) to learn this prompt and improve future trajectory prediction. We call our learnable prompt a latent corridor: a non-physical corridor that guides human behavior in this scene.
@@
In our experiments, learning latent corridors for prediction adaptation enables up to $23.9\%$ prediction accuracy improvement on simulated data from MOTSynth [12] compared to a non-adaptive predictor that takes as input the scene and has seen all the same data. On real data from MOT and WildTrack [7], we see a $16.4\%$ improvement on non-adaptive predictors and a $11.2\%$ improvement with adapting via our method as opposed to just fine-tuning. We find that our latent corridor adaptation can improve prediction performance even with a very small amount of data (e.g. a couple of humans for a few seconds), and continually improves as more data (e.g., many humans for a few minutes) is observed over time. Qualitatively, we observe that learning latent corridors can inject 3D ground plane awareness into prediction performed from an arbitrary camera view angle, can help the predictor adapt to obstacles or occlusions in the scene, and learn scene-specific patterns of human behavior.
@@
    We demonstrate qualitative and quantitative improvements of up to $23.9\%$ ADE improvement in MOTSynth simulated data and $16.4\%$ ADE in MOT and Wildtrack real pedestrian data.
@@ 2 Related Work
Human Trajectory Prediction is a well-studied problem with a long and rich history [33]. Prediction methods started from simple models such as social forces [5, 15], and later added multimodality through Gaussian processes [38]. Recently, modern learning-based predictors have made significant progress in incorporating social interactions between humans as modelled via RNNs [40, 16, 35], GANs [13], or conditional VAEs [27], all while generating multimodal future trajectories. Another line of works incorporate the scene context in predictions, getting neural network features from a scene map [20, 39, 46, 22, 34], and more recently using transformers [30, 36]. A few works have incorporated the scene context by projecting the trajectories into heatmaps so that the network can reason in image space [28, 6], allowing for longer-term predictions. Our work builds upon these advances by adapting an RGB scene-aware pre-trained trajectory predictor [28] over time.
Adapting Human Trajectory Predictors. In the past few years, there has been a growing interest in lightweight ways to modify pre-trained predictors to new data. The key differences between these works lie in 1) what they are adapting to, and 2) how they adapt. Several works focus on cross-domain transfer, wherein a predictor trained for trajectory prediction in domain A (e.g., New York) is adapted to work in domain B (e.g., London). These works leverage architectures that partition generic trajectory prediction from domain-specific features [45, 42], leverage adaptive meta-learning via Kalman flitering [18], or simply finetune the prediction network on new data [25]. Other works focus on online adaptation over time, for example by adapting to different agent dynamics using recursive least squares [1, 10].  [47] carries out continual learning on 2D birds-eye view scenes for pedestrian trajectory prediction, using a memory bank to store embedded pairs of trajectory pasts and futures. In this work, we focus on adaptation over time within a specific deployment domain, but propose a novel prompting-based adaptation method which learns scene-specific information in a scalable way from in-the-wild data.
@@ 3 Problem Formulation
Here, we present a formalization of our proposed adpative trajectory prediction problem. We seek to generate accurate future trajectories of $N$ agents in a scene. At any time $\tau$, let $\mathbf{o}_{\tau}:=o_{\tau-H:\tau}\in\mathcal{O}$ be the $H$-step history of observations input into the predictor, $\mathbf{y}_{\tau}:=y_{\tau+1:\tau+T}\in\mathcal{Y}$ be the ground-truth $T$-step future trajectory, and $\hat{\mathbf{y}}_{\tau}:=\hat{y}_{\tau+1:\tau+T}\in\mathcal{Y}$ be the prediction outputs. We assume access to a base predictor $\mathcal{P}$ pre-trained on a human motion dataset consisting of past and ground-truth future trajectories, $\mathcal{D}=\{\mathbf{o}_{i},\mathbf{y}_{i}\}^{K}_{i=1}$ to minimize a loss function on trajectories, $\ell$ (e.g., mean squared error).
Adaptive Trajectory Prediction (ATP). We aim to adapt $\mathcal{P}$ to a deployment scene by observing new human trajectory data over a time interval. Let the dataset of new human trajectory data be $\mathcal{D}^{\prime}_{0:T}$ collected over the $T$ step time interval. The deployment scenes being adapted to in $\mathcal{D}^{\prime}$ can be real or synthetic, can be of a new physical environment or same as the pre-training dataset, and can be obtained by a new or previously seen camera configuration. However, we assume that human pedestrians are present in the scene and they are navigating an outdoor environment. We adapt to scenes captured for a similar amount of time, between $T=45$ seconds and $5$ minutes. We aim to adapt to scene-specific characteristics and fluctuating events beyond those that can be exploited by conditioning only on a segmentation map.
@@
Intuitively, as the value of $t$ increases and the adapted predictor sees more data, the performance of $\mathcal{A}[\mathcal{P}]$ should improve over the pre-trained predictor $\mathcal{P}$. At deployment, $\mathcal{A}[\mathcal{P}]$ should have learned scene context-specific characteristics of human behavior.
@@ 4 Adaptation via Learned Latent Corridors
To adapt our trajectory predictor efficiently, we take inspiration from language-based prompt-tuning [23] and instantiate it within our adaptive trajectory prediction problem. Specifically, we augment a frozen base predictor with a learnable latent prompt we call a latent corridor, a non-physical corridor that informs the predictor of human behavior patterns in a specific deployment scene. In this section, we detail our predictor architecture, prompt representation and training, and investigate a suite of prompting configurations that are compatible with latent corridors.
@@ 4.2 Representing & Learning Latent Corridors
Our key idea for adapting trajectory predictors is to augment the frozen base model $\mathcal{P}$ with a set of trainable prompts, called latent corridors, which learn new trends in how humans interact with the scene or with each other. For effective adaptation, we seek two key properties for our latent corridors: parameter-efficient (i.e., we want to minimize the number of parameters that are tuned from new deployment data) and spatially scene-grounded (i.e., the latent should have pixel-wise alignment with the scene). We first outline our prompting approach, and then discuss a compact but spatially-grounded representation of the prompt that is amenable to parameter-efficient learning.
Latent corridor prompt. For each of the $K$ deployment scenes, we introduce a unique trainable prompt, $p_{k}\in\mathbb{R}^{h\times w}$, of the same spatial size as the image $I$ that is input into the predictor (left, Fig. 2). For a network that has adapted to $K$ scenes, we have a set $K$ of prompts $p:=\{p_{0},p_{1},\ldots,p_{K}\}$. Thus, our adaptation rule is
$\mathcal{A}:=\mathbf{M}_{t}\oplus p\quad\forall t\in\{\tau-H,\ldots,\tau\}.$ (3)
The prompt is summed element-wise to each of the input heatmaps corresponding to the observed trajectories; let $\widetilde{\mathbf{M}}=\mathbf{M}\oplus p$ denote this for any original $\mathbf{M}$. See Sec. 8 for an ablation on prompt location. The adapted predictor takes as input $\mathcal{P}([\widetilde{\mathbf{M}}_{\tau-H:\tau},S])$.
A compact but spatially-grounded representation. In Equation (3), the prompt is assumed to be the same size as the image, $h\times w$, which is nice for spatial alignment between the prompt and scene. However, in our experiments, where all images are of size $h=288$ and $w=480$, this naive image-based representation requires learning an additional $138$K parameters. Considering that our base predictor $\mathcal{P}$ has $\sim 900$K trainable parameters, this prompt increases the model parameter size by over $15\%$ and makes it infeasible to learn meaningful latent corridors from small amounts of new human trajectories (see Sec. 8). Instead, we propose to represent the prompt as a low-rank representation with rank $1$. We initialize our latent corridor as a vector of dimension $h+w$ using Kaiming initialization [14], and parameterize the full prompt as the outer product of the $h$ and $w$ dimensional vectors. This lightweight representation preserves the spatial relationship between the prompt, scene, and trajectory heatmaps and is significantly more compact: it increases the model parameter size by less than $0.1\%$.
@@ 4.3 Prompting configurations
    Per-scene finetuning (LC + Per-Scene FT): If one seeks to maximize performance within a specific deployment scene, one can jointly finetune $K$ predictor heads $\mathcal{P}_{P_{k}}$ with $K$ latent corridors per scene. While this results in the need for a unique predictor for each deployment scene, we find empirically that this method achieves best in-scene adaptation performance.

### v2

@@ Abstract
Human trajectory prediction is typically posed as a zero-shot generalization problem: a predictor is learnt on a dataset of human motion in training scenes, and then deployed on unseen test scenes. While this paradigm has yielded tremendous progress, it fundamentally assumes that trends in human behavior within the deployment scene are constant over time. As such, current prediction models are unable to adapt to transient human behaviors, such as crowds temporarily gathering to see buskers, pedestrians hurrying through the rain and avoiding puddles, or a protest breaking out. We formalize the problem of context-specific adaptive trajectory prediction and propose a new adaptation approach inspired by prompt tuning called latent corridors. By augmenting the input of a pre-trained human trajectory predictor with learnable image prompts, the predictor improves in the deployment scene by inferring trends from extremely small amounts of new data (e.g., 2 humans observed for 30 seconds). With less than $0.1\%$ additional model parameters, we see up to $23.9\%$ ADE improvement in MOTSynth simulated data and $16.4\%$ ADE in MOT and Wildtrack real pedestrian data. Qualitatively, we observe that latent corridors imbue predictors with an awareness of scene geometry and context-specific human behaviors that non-adaptive predictors struggle to capture.
Keywords:
human trajectory prediction, adaptation, test-time training, image prompt tuning
@@ 1 Introduction
Figure 1: Adaptive trajectory prediction. (left) Given a history of human behavior (shown in black), the pre-trained predictor $\mathcal{P}$ is unable to understand deployment scene-specific behavior trends, like people entering a subterranean subway entrance (bottom row) or mostly choosing to traverse the staircase as opposed to exploring other parts of the scene at nighttime (top row). (right) When adapting, the number of people and amount of time determine the total number of trajectories observed, and we denote this time-dependent quantity human-seconds. Here, the three columns correspond to our method trained for a very small (left), medium (middle), and large amount of human-seconds (right). Our adaptive latent corridors approach enables $\mathcal{P}$ to quickly learn context-specific trends, improving predictions with even small amounts of data, and closing the gap between the ground-truth (green) and predicted behavior (orange). For example, in the middle row, $\mathcal{P}$ predicts the person will move towards the camera, but as our method sees more human-seconds of data, it adapts to the trend that at this point of scene capture in the plaza, people tend to avoid the center of the plaza and instead move diagonally across it.
@@
Human motion prediction is a fundamental skill for intelligent systems to effectively navigate the world, assist end-users, and visually survey a scene. To date, learning-based human trajectory prediction has been extensively studied, making huge strides in predicting multimodal future behavior [28], modeling multi-agent social interactions [2, 18, 40], and accounting for scene context [56, 27]. Some approaches have exploited the fact that humans tend to move in similar patterns by storing prior trajectories in a bank and matching a current observed trajectory to the closest stored one [31, 30, 50, 54].
@@
However, a fundamental assumption underlies the current trajectory prediction paradigm: predictors are assumed to be deployed within an unchanging context. Nevertheless, in the real world, the context—and thus patterns of human behavior—will inevitably change over time, as Whyte observed in his famous study of urban public spaces [49]. For example, a new subway entrance being built causes people to descend and exit in new patterns (bottom row, Fig. 1), and icy sidewalks in the winter will be traversed differently than clear ones in the summer. These changes happen on shorter temporal horizons, too: Humans rushing to work during the day will largely ignore each other while nighttime party-goers will walk in cliques. A college campus will be quiet while class is in session and chaotic in the 10 minute period where students rush between classes. Even if data is captured from exactly the same location, human behavior patterns can change in minutes or hours. When faced with such changes, the performance of existing trajectory predictors immediately degrades.
@@
Here, we study how to efficiently adapt pre-trained trajectory predictors to observed human behavior within a deployment scene (right, Fig. 1). Our technical approach takes inspiration from the recently popularized paradigm of prompt-tuning in large language models [23], but instantiates this idea within the adaptive trajectory prediction problem. Specifically, we augment the input to an existing human trajectory predictor with a set of latent image prompts, one per each instance of a deployment scene we wish to adapt to. We leave all or most of the predictor frozen, and tune each input prompt with the same trajectory prediction loss that is used to train the base predictor, letting cues in human behaviour be captured while preserving useful structure in the original predictor (e.g., even if a new subway entrance lets people descend into the ground, other people may still move towards the existing above-ground coffee shops). Our adaptation is data-efficient: even extremely small amounts of new human trajectories (e.g., $<1$ minute of human seconds, corresponding to 2 humans observed over 30 seconds) can leak sufficient information about how humans interact with the scene (e.g., new subway entrance) or with each other (e.g., nighttime social cliques) to learn this prompt and improve future trajectory prediction. We call our learnable prompt a latent corridor: a non-physical corridor that guides human behavior in this scene.
Given the small amounts of data, latent corridors train quickly—on average within minutes on a single 2080 Ti GPU—even without optimizations to improve training speed. With more compute resources and training optimization, latent corridors hold the potential to learn in real time, adapting to transient events such a fire starting or a musician busking. Even with minimal compute and optimization, we also find useful offline applications of latent corridors for adaptation to repeated events. For example, imagine analyzing pedestrian flow in a downtown business district — it will differ during the rush hour before work on the weekday and on a weekend morning. During the course of a week, latent corridors could be trained to adapt to the varying scene and context specific behaviours present on different days, times, and surveillance camera viewpoints. Once training is complete, adapted models could be used online for months, until an event such as construction starting or a change of season causes a significant change in behaviour. Our method is lightweight, making swapping latent corridors for different times of day or locations inexpensive, even on device.
@@
In our experiments, learning latent corridors for prediction adaptation enables $23.9\%$ prediction accuracy improvement on simulated data from MOTSynth [12] compared to a non-adaptive predictor that takes as input the scene and has seen all the same training data. On real data from MOT and WildTrack [7], we see a $16.4\%$ improvement on non-adaptive predictors and an $11.2\%$ improvement with adapting via our method as opposed to just fine-tuning, and on in-the-wild webcam data we see respective improvements of $26.8\%$ and $20.0\%$. Latent corridor adaptation can improve prediction performance even with a very small amount of data, and continually improves as more data is observed over time.
@@
    We demonstrate qualitative and quantitative improvements of up to $23.9\%$ ADE improvement in MOTSynth simulated data, $16.4\%$ ADE in MOT and Wildtrack real pedestrian data, and $26.8\%$ ADE on in-the-wild webcam data, over a scene-aware baseline.
@@ 2 Related Work
Human Trajectory Prediction is a well-studied problem with a long and rich history [37]. Prediction methods started from simple models such as social forces [5, 15], and later added multimodality through Gaussian processes [43]. Recently, modern learning-based predictors have made significant progress in incorporating social interactions between humans modelled via RNNs [45, 17, 40], GANs [13], or conditional VAEs [28], all while generating multimodal future trajectories. Another line of work incorporates the scene context in predictions, getting neural network features from a scene map [20, 44, 52, 22, 38, 30, 29], embedding gridded scene context in an LSTM [29, 53], and more recently using transformers [33, 39]. A few works incorporate scene context by projecting the trajectories into heatmaps so that the network can reason in image space [27, 6], allowing for longer-term predictions. Our work builds upon these advances by adapting an RGB scene-aware pre-trained trajectory predictor [27] over time.
Adapting Human Trajectory Predictors. In the past few years, there has been a growing interest in lightweight ways to modify pre-trained predictors to new data. The key differences between these works lie in 1) what they are adapting to, and 2) how they adapt. Several works focus on cross-domain transfer, wherein a predictor trained for trajectory prediction in domain A (e.g., New York) is adapted to work in domain B (e.g., London). These works leverage architectures that partition generic trajectory prediction from domain-specific features [51, 47], leverage adaptive meta-learning via Kalman flitering [16], or simply finetune the prediction network on new data [25]. Other works focus on online adaptation over time, for example by adapting to different agent dynamics using recursive least squares [1, 10]. A few approaches carry out continual learning by storing embedded past/future trajectory pairs in a memory bank [30, 50, 54] or clustering them [42], predicting by matching an individuals history to the most similar stored past. Going a step further, [31] instead stores representative group trajectories in a scene and during inference, refines the selected trajectory with scene segmentation. In contrast, our novel prompting-based adaptation approach is memory-efficient, using constant extra parameter space regardless of trajectory dataset size, and inherently utilizes the scene segmentation, not requiring extra refinement steps. It also directly augments an existing model as opposed to requiring a new, specialized architecture.
@@ 3 Problem Formulation
Here, we present a formalization of our proposed adaptive trajectory prediction problem. We seek to generate accurate future trajectories of $N$ agents in a scene. At any time $\tau$, let $\mathbf{o}_{\tau}:=o_{\tau-H:\tau}\in\mathcal{O}$ be the $H$-step history of observations input into the predictor, $\mathbf{y}_{\tau}:=y_{\tau+1:\tau+T}\in\mathcal{Y}$ be the ground-truth $T$-step future trajectory, and $\hat{\mathbf{y}}_{\tau}:=\hat{y}_{\tau+1:\tau+T}\in\mathcal{Y}$ be the prediction outputs. We assume access to a base predictor $\mathcal{P}$ pre-trained on a human motion dataset consisting of past and ground-truth future trajectories, $\mathcal{D}=\{\mathbf{o}_{i},\mathbf{y}_{i}\}^{K}_{i=1}$ to minimize a loss function on trajectories, $\ell$ (e.g., mean squared error).
Adaptive Trajectory Prediction (ATP). We aim to adapt $\mathcal{P}$ to a deployment scene by observing new human trajectory data over a time interval. Let the dataset of new human trajectory data be $\mathcal{D}^{\prime}_{0:T}$ collected over the $T$ step time interval. Here, $\mathcal{D}^{\prime}$ can contain data from real or synthetic deployment scenes, a new physical environment or one from the pre-training dataset, a new or previously seen time period (e.g., nightime vs. daytime), and can be obtained by a new or previously seen camera pose. However, we assume that human pedestrians are present in the scene and they are navigating an outdoor environment. We adapt to scenes captured for a similar amount of time, between $T=45$ seconds and $5$ minutes. We aim to adapt to scene-specific characteristics and fluctuating events beyond those that can be exploited by conditioning only on a segmentation map.
@@
Intuitively, as the value of $t$ increases and the adapted predictor sees more data, as long as the context within the window $[0:T]$ is consistent, the performance of $\mathcal{A}[\mathcal{P}]$ should improve over the pre-trained predictor $\mathcal{P}$. At deployment, $\mathcal{A}[\mathcal{P}]$ should have learned scene context-specific characteristics of human behavior.
@@ 4 Adaptation via Learned Latent Corridors
To adapt our trajectory predictor efficiently, we take inspiration from language-based prompt-tuning [23] and instantiate it within our adaptive trajectory prediction problem. Specifically, we augment a frozen base predictor with a learnable latent prompt we call a latent corridor, a non-physical corridor that informs the predictor of human behavior patterns in a specific deployment scene. In this section, we detail our predictor architecture, prompt representation and training, and investigate a suite of latent corridor prompting configurations.
@@ 4.2 Representing & Learning Latent Corridors
Our key idea for adapting trajectory predictors is to augment the frozen base model $\mathcal{P}$ with a set of trainable prompts, called latent corridors, which learn new trends in how humans interact with the scene or with each other. For effective adaptation, we seek two key properties for our latent corridors: parameter-efficient (i.e., we want to minimize the number of parameters that are tuned from new deployment data) and spatially scene-grounded (i.e., the latent should have pixel-wise alignment with the scene image). We first outline our prompting approach, and then discuss a compact but spatially-grounded representation of the prompt that is amenable to parameter-efficient learning.
Latent corridor prompt. For each of the $K$ deployment scenes, we introduce a unique trainable prompt, $p_{k}\in\mathbb{R}^{h\times w}$, of the same spatial size as the image $I$ that is input into the predictor (left, Fig. 2). For a network that has adapted to $K$ scenes, we have a set $K$ of prompts $p:=\{p_{0},p_{1},\ldots,p_{K}\}$. Thus, our adaptation rule to scene $k$ is
$\mathcal{A}:=\mathbf{M}_{t}\oplus p_{k}\quad\forall t\in\{\tau-H,\ldots,\tau\}.$ (3)
The prompt is summed element-wise to each of the input heatmaps corresponding to the observed trajectories; let $\widetilde{\mathbf{M}}=\mathbf{M}\oplus p$ denote this for any original $\mathbf{M}$. See Supplement Sec.1 for an ablation on prompt location. The adapted predictor takes as input $\mathcal{P}([\widetilde{\mathbf{M}}_{\tau-H:\tau},S])$.
A compact but spatially-grounded representation. In Equation (3), the prompt is assumed to be the same size as the image, $h\times w$, which is nice for spatial alignment between the prompt and scene. However, in our experiments, where all images are of size $h=288$ and $w=480$, this naive image-based representation requires learning an additional $138$K parameters. Considering that our base predictor $\mathcal{P}$ has $\sim 900$K trainable parameters, this prompt increases the model parameter size by over $15\%$. Instead, we propose a low-rank representation of the prompt with rank $1$. We initialize our latent corridor as a vector of dimension $h+w$ using Kaiming initialization [14], and parameterize the full prompt as the outer product of the $h$ and $w$ dimensional vectors. This lightweight representation preserves the spatial relationship between the prompt, scene, and trajectory heatmaps and is significantly more compact: it increases the model parameter size by less than $0.1\%$. We find empirically that this rank 1 matrix performs similarly to a full rank matrix.
@@ 4.3 Prompting configurations
    Per-scene finetuning (LC + Per-Scene FT): If one seeks to maximize performance within a specific deployment scene, one can jointly finetune $K$ predictor heads $\mathcal{P}_{P_{k}}$ with $K$ latent corridors (one per scene). While this results in the need for a unique predictor for each deployment scene, we find empirically that this method achieves best in-scene adaptation performance.